OpenStudy (anonymous):

Simplify the expression twenty-one minus seven times two plus nine. Show your work. Then, show how the solution is affected if addition and subtraction are not applied from left to right.

OpenStudy (j2lie):

21-7+2=16(from left to right). 21-7+2=12(not from left to right) The answer decreased if you don't do it from left to right.

OpenStudy (anonymous):